A wheel rotates with a constant angular velocity of 10 rad per second.

A, compute the Radial acceleration of a point 0.5m from the axis

B, find the tangential velocity at that point


1
Expert's answer
2019-12-16T11:11:11-0500

The radial acceleration of a point:

"a=\omega^2r=(10)^2(0.5)=50\frac{m}{s^2}"

The  tangential velocity at that point:

"v=\omega r=(10) (0.5)=5\frac{m}{s }"